# Multitype entanglement dynamics induced by exceptional points

### v1

@@ C.2 First-order degenerate and non-degenerate perturbation theory with Ising type interaction
The perturbation matrix in the subspace spanned by $\left\{\left|{{\widetilde{\psi}}_{+-}}\right\rangle,\left|{{\widetilde{\psi}}_{-+}}\right\rangle\right\}$ and their adjoint states $\left\{\left\langle{{\widetilde{\overline{\psi}}}_{+-}}\right|,\left\langle{{\widetilde{\overline{\psi}}}_{-+}}\right|\right\}$. This matrix can be obtained as
@@
Here, $J=-2\xi$. Noted that ${{H}_{\operatorname{int}}}$ is Hermitian and this submatrix is real with eigenvalues 0 and $-\frac{2J{{\gamma}^{2}}}{{{\eta}^{2}}}$. The eigenstates read as $\left(\left|{{\psi}_{-+}}\right\rangle\mp\left|{{\psi}_{+-}}\right\rangle\right)/\sqrt{2}$, which means that we can choose a new basis for the degenerate subspace, specifically,
@@
Noted that this type of interaction Hamiltonian ${{H}_{\operatorname{int}}}$ is diagonal and the state $\left|{{\widetilde{\psi}}_{1}}\right\rangle$is equal to $\left|{{\widetilde{\overline{\psi}}}_{1}}\right\rangle$. The corresponding eigenvalues are given by
@@
It is easy to find that $\left|{{\widetilde{\psi}}_{1}}\right\rangle=\left|{{\widetilde{\overline{\psi}}}_{1}}\right\rangle$. Applying to the first-order non-degenerate perturbation theory to our system of two weakly coupled non-Hermitian spins. The perturbed eigenvalues are given by
$\Lambda_{1}$ $=\lambda_{1}+0=-\frac{i\gamma}{2}-2\xi,$ (35)
$\overline{\Lambda}_{1}$ $=\overline{\lambda}_{1}+0=\frac{i\gamma}{2}-2\xi,$
$\Lambda_{2}$ $=\lambda_{2}-\frac{2J\gamma^{2}}{\eta^{2}}=-\frac{i\gamma}{2}-2\xi-\frac{2J\gamma^{2}}{\eta^{2}},$
$\overline{\Lambda}_{2}$ $=\overline{\lambda}_{2}-\frac{2J\gamma^{2}}{\eta^{2}}=\frac{i\gamma}{2}-2\xi-\frac{2J\gamma^{2}}{\eta^{2}}.$
@@
Finding that the given interaction Hamiltonian is already diagonal in the basis $\left\{\left|{{\widetilde{\psi}}_{1}}\right\rangle,\left|{{\widetilde{\psi}}_{2}}\right\rangle\right\}$ of the Hilbert space and $\left\{\left|{{\widetilde{\overline{\psi}}}_{1}}\right\rangle,\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle\right\}$ of the dual space. Therefore, we need to consider the first-order unperturbed eigenstates of the non-degenerate subspace.
@@
$\left|{{\Psi}_{1}}\right\rangle=\left|{{\overline{\Psi}}_{1}}\right\rangle$ $=\left|{{\widetilde{\psi}}_{1}}\right\rangle+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{--}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{1}}\right\rangle}{{{\lambda}_{1}}-{{\lambda}_{--}}}\left|{{\widetilde{\psi}}_{--}}\right\rangle$ (36)
@@
$+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{++}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{1}}\right\rangle}{{{\lambda}_{1}}-{{\lambda}_{++}}}\left|{{\widetilde{\psi}}_{++}}\right\rangle$
$=\frac{1}{\sqrt{2}}\left(\begin{matrix}0\\ -1\\ 1\\ 0\\ \end{matrix}\right),$
@@
$\left|{{\Psi}_{2}}\right\rangle$ $=\left|{{\widetilde{\psi}}_{2}}\right\rangle+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{--}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{--}}}\left|{{\widetilde{\psi}}_{--}}\right\rangle$ (37)
$+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{++}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{++}}}\left|{{\widetilde{\psi}}_{++}}\right\rangle$
$=\frac{1}{\sqrt{2}{{\eta}^{3}}}\left(\begin{matrix}-4\Omega{{\eta}^{2}}-i16\gamma J\Omega\\ i\gamma{{\eta}^{2}}\\ i\gamma{{\eta}^{2}}\\ 4\Omega{{\eta}^{2}}-i16\gamma J\Omega\\ \end{matrix}\right),$
$\left|{{\overline{\Psi}}_{2}}\right\rangle$ $=\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle+\frac{\left\langle{{\widetilde{\psi}}_{--}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{--}}}\left|{{\widetilde{\overline{\psi}}}_{--}}\right\rangle$ (38)
$+\frac{\left\langle{{\widetilde{\psi}}_{++}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{++}}}\left|{{\widetilde{\overline{\psi}}}_{++}}\right\rangle$
$=\frac{1}{\sqrt{2}{{\eta}^{3}}}\left(\begin{matrix}-4\Omega{{\eta}^{2}}+i16\gamma J\Omega\\ -i\gamma{{\eta}^{2}}\\ -i\gamma{{\eta}^{2}}\\ 4\Omega{{\eta}^{2}}+i16\gamma J\Omega\\ \end{matrix}\right).$
@@
The above perturbation theory gives the basis states, i.e., $\left|{{\psi}_{++}}\right\rangle,\left|{{\psi}_{--}}\right\rangle,\left|{{\psi}_{1}}\right\rangle,\left|{{\psi}_{2}}\right\rangle$ in the Hibert space $H$ as well as the corresponding adjoint states $\left|{{\overline{\psi}}_{++}}\right\rangle,\left|{{\overline{\psi}}_{--}}\right\rangle,\left|{{\overline{\psi}}_{1}}\right\rangle,\left|{{\overline{\psi}}_{2}}\right\rangle$ in the dual space ${{H}^{*}}$. It is worth to note that the above basis states are both normalized
@@
$\Lambda_{--}$ $=\lambda_{--}+\left\langle\widetilde{\overline{\psi}}_{--}\middle|H_{\text{int}}\middle|\widetilde{\psi}_{--}\right\rangle$ (44)
$=\frac{1}{2}\left(-i\gamma-4\xi-\eta\right)+\frac{16J\Omega^{2}}{{\eta^{2}}},$
$\Lambda_{++}$ $=\lambda_{++}+\left\langle\widetilde{\overline{\psi}}_{++}\middle|H_{\text{int}}\middle|\widetilde{\psi}_{++}\right\rangle$
$=\frac{1}{2}\left(-i\gamma-4\xi+\eta\right)+\frac{16J\Omega^{2}}{{\eta^{2}}}.$
@@ C.3 Entanglement between two coupled non-Hermitian qubits
$\left|\psi(t)\right\rangle$ $=U\left|\psi(0)\right\rangle=\left\langle\overline{\Psi}_{++}\middle|\psi(0)\right\rangle e^{-it\Lambda_{++}}\left|\Psi_{++}\right\rangle\quad+\left\langle\overline{\Psi}_{--}\middle|\psi(0)\right\rangle e^{-it\Lambda_{--}}\left|\Psi_{--}\right\rangle$ (45)
$\quad+\left\langle\overline{\Psi}_{1}\middle|\psi(0)\right\rangle e^{-it\Lambda_{1}}\left|\Psi_{1}\right\rangle\quad+\left\langle\overline{\Psi}_{2}\middle|\psi(0)\right\rangle e^{-it\Lambda_{2}}\left|\Psi_{2}\right\rangle$
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) (Grants No. 62075004, No. 11804018, No. U23A20481, No. 62275010, No. 12074027 and No. 12104252) and Beijing Municipal Natural Science Foundation (Grants No. 4212051 and No. 1232027).

### v2

@@ C.2 First-order degenerate and non-degenerate perturbation theory with Ising type interaction
The perturbation matrix in the subspace is spanned by $\left\{\left|{{\widetilde{\psi}}_{+-}}\right\rangle,\left|{{\widetilde{\psi}}_{-+}}\right\rangle\right\}$ and their adjoint states $\left\{\left\langle{{\widetilde{\overline{\psi}}}_{+-}}\right|,\left\langle{{\widetilde{\overline{\psi}}}_{-+}}\right|\right\}$. This matrix can be obtained as
@@
Here, $J=-2\xi$. Note that ${{H}_{\operatorname{int}}}$ is Hermitian and this submatrix is real with eigenvalues 0 and $-\frac{2J{{\gamma}^{2}}}{{{\eta}^{2}}}$. The eigenstates read as $\left(\left|{{\psi}_{-+}}\right\rangle\mp\left|{{\psi}_{+-}}\right\rangle\right)/\sqrt{2}$, which means that we can choose a new basis for the degenerate subspace, specifically,
@@
Note that this type of interaction Hamiltonian ${{H}_{\operatorname{int}}}$ is diagonal and the state $\left|{{\widetilde{\psi}}_{1}}\right\rangle$is equal to $\left|{{\widetilde{\overline{\psi}}}_{1}}\right\rangle$. The corresponding eigenvalues are given by
@@
It is easy to find that $\left|{{\widetilde{\psi}}_{1}}\right\rangle=\left|{{\widetilde{\overline{\psi}}}_{1}}\right\rangle$. We apply the first-order non degenerate perturbation theory to our system of two weakly coupled non-Hermitian spins. The perturbed eigenvalues are given by
$\Lambda_{1}$ $=\lambda_{1}+0=-\frac{i\gamma}{2}-2\xi,\quad\overline{\Lambda}_{1}=\overline{\lambda}_{1}+0=\frac{i\gamma}{2}-2\xi,$ (39)
$\Lambda_{2}$ $=\lambda_{2}-\frac{2J\gamma^{2}}{\eta^{2}}=-\frac{i\gamma}{2}-2\xi-\frac{2J\gamma^{2}}{\eta^{2}},\quad\overline{\Lambda}_{2}=\overline{\lambda}_{2}-\frac{2J\gamma^{2}}{\eta^{2}}=\frac{i\gamma}{2}-2\xi-\frac{2J\gamma^{2}}{\eta^{2}}.$
@@
We find that the given interaction Hamiltonian is already diagonal in the basis $\left\{\left|{{\widetilde{\psi}}_{1}}\right\rangle,\left|{{\widetilde{\psi}}_{2}}\right\rangle\right\}$ of the Hilbert space and $\left\{\left|{{\widetilde{\overline{\psi}}}_{1}}\right\rangle,\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle\right\}$ of the dual space. Therefore, we need to consider the first-order unperturbed eigenstates of the non-degenerate subspace.
@@
$\left|{{\Psi}_{1}}\right\rangle=\left|{{\overline{\Psi}}_{1}}\right\rangle=\left|{{\widetilde{\psi}}_{1}}\right\rangle+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{--}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{1}}\right\rangle}{{{\lambda}_{1}}-{{\lambda}_{--}}}\left|{{\widetilde{\psi}}_{--}}\right\rangle+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{++}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{1}}\right\rangle}{{{\lambda}_{1}}-{{\lambda}_{++}}}\left|{{\widetilde{\psi}}_{++}}\right\rangle=\frac{1}{\sqrt{2}}{{\left(\begin{matrix}0&-1&1&0\\ \end{matrix}\right)}^{\text{T}}},$ (40)
@@
$\left|{{\Psi}_{2}}\right\rangle=\left|{{\widetilde{\psi}}_{2}}\right\rangle+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{--}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{--}}}\left|{{\widetilde{\psi}}_{--}}\right\rangle+\frac{\left\langle{{\widetilde{\overline{\psi}}}_{++}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\psi}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{++}}}\left|{{\widetilde{\psi}}_{++}}\right\rangle=\frac{1}{\sqrt{2}{{\eta}^{3}}}\left(\begin{matrix}-4\Omega{{\eta}^{2}}-i16\gamma J\Omega\\ i\gamma{{\eta}^{2}}\\ i\gamma{{\eta}^{2}}\\ 4\Omega{{\eta}^{2}}-i16\gamma J\Omega\\ \end{matrix}\right),$ (41)
@@
$\left|{{\overline{\Psi}}_{2}}\right\rangle=\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle+\frac{\left\langle{{\widetilde{\psi}}_{--}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{--}}}\left|{{\widetilde{\overline{\psi}}}_{--}}\right\rangle+\frac{\left\langle{{\widetilde{\psi}}_{++}}\right|{{H}_{\operatorname{int}}}\left|{{\widetilde{\overline{\psi}}}_{2}}\right\rangle}{{{\lambda}_{2}}-{{\lambda}_{++}}}\left|{{\widetilde{\overline{\psi}}}_{++}}\right\rangle=\frac{1}{\sqrt{2}{{\eta}^{3}}}\left(\begin{matrix}-4\Omega{{\eta}^{2}}+i16\gamma J\Omega\\ -i\gamma{{\eta}^{2}}\\ -i\gamma{{\eta}^{2}}\\ 4\Omega{{\eta}^{2}}+i16\gamma J\Omega\\ \end{matrix}\right).$ (42)
@@
The above perturbation theory gives the basis states, i.e., $\left|{{\psi}_{++}}\right\rangle,\left|{{\psi}_{--}}\right\rangle,\left|{{\psi}_{1}}\right\rangle,\left|{{\psi}_{2}}\right\rangle$ in the Hilbert space $H$ as well as the corresponding adjoint states $\left|{{\overline{\psi}}_{++}}\right\rangle,\left|{{\overline{\psi}}_{--}}\right\rangle,\left|{{\overline{\psi}}_{1}}\right\rangle,\left|{{\overline{\psi}}_{2}}\right\rangle$ in the dual space ${{H}^{*}}$. It is worth to note that the above basis states are both normalized
@@
$\Lambda_{--}=\lambda_{--}+\left\langle\widetilde{\overline{\psi}}_{--}\middle|H_{\text{int}}\middle|\widetilde{\psi}_{--}\right\rangle=\frac{1}{2}\left(-i\gamma-4\xi-\eta\right)+\frac{16J\Omega^{2}}{{\eta^{2}}},$ (48)
$\Lambda_{++}=\lambda_{++}+\left\langle\widetilde{\overline{\psi}}_{++}\middle|H_{\text{int}}\middle|\widetilde{\psi}_{++}\right\rangle=\frac{1}{2}\left(-i\gamma-4\xi+\eta\right)+\frac{16J\Omega^{2}}{{\eta^{2}}}.$
@@ C.3 Entanglement between two coupled non-Hermitian qubits
$\left|\psi(t)\right\rangle$ $=U\left|\psi(0)\right\rangle$ (49)
$=\big{\langle}\overline{\Psi}_{++}\big{|}\psi(0)\big{\rangle}e^{-it\Lambda_{++}}\left|\Psi_{++}\right\rangle+\big{\langle}\overline{\Psi}_{--}\big{|}\psi(0)\big{\rangle}e^{-it\Lambda_{--}}\left|\Psi_{--}\right\rangle$
$\quad+\big{\langle}\overline{\Psi}_{1}\big{|}\psi(0)\big{\rangle}e^{-it\Lambda_{1}}\left|\Psi_{1}\right\rangle+\big{\langle}\overline{\Psi}_{2}\big{|}\psi(0)\big{\rangle}e^{-it\Lambda_{2}}\left|\Psi_{2}\right\rangle$
@@
Figure C2: Time evolution of the modulus $\left|j\right|(j=\alpha,\beta,\varsigma,\delta)$ for each of the four complex amplitudes of the two-qubit state $\left|\widetilde{\psi}\right\rangle=\alpha\left|aa\right\rangle+\beta\left|ab\right\rangle+\zeta\left|ba\right\rangle+\delta\left|bb\right\rangle$ for $\xi=0$ (a) and $\xi/\gamma=0.006$ (b-d). In the absence of Ising interaction, the evolution of $\left|j\right|$ is focused in the $\mathcal{P}\mathcal{T}$-symmetric phase. In the presence of Ising interaction, the evolution of $\left|j\right|$ has been shown in three different regions, i.e., in the $\mathcal{P}\mathcal{T}$-symmetry-broken phase (b), the mixed phase (c) and $\mathcal{P}\mathcal{T}$-symmetric phase (d), respectively. The given driving amplitude are chosen as $\Omega/\gamma=0.23$ (b), $\Omega/\gamma=0.28$ (c) and $\Omega/\gamma=0.3$ (d).
According to Eq. (52), the multitype entanglement dynamics (type I, II, III) can also be described in the aspect of the population of the basis state. Corresponding to the population at each basis state they exhibits different oscillatory behaviors in the three regions divided by the original EP4 and EP2 (as shown in Fig. C2), and different entanglement dynamics can be found in the three regions, accordingly. Compared with the population evolutions of the basis states $\left|bb\right\rangle$ and $\left|aa\right\rangle$ in the three different regions, we can find that the distorted Rabi-like oscillation of the population of $\left|bb\right\rangle$ and $\left|aa\right\rangle$ in the PTS, as the black and pink curves shown in Fig. C2(a) and  C2(d) regardless of whether there exists interaction. Moreover, the population evolution of the states $\left|ab\right\rangle$ and $\left|ba\right\rangle$ monotonically increase during the system evolving to a steady state in the PI [Fig. C2(b)] and the mixed phase [Fig. C2(c)]. Strikingly, the populations evolution of the basis states $\left|bb\right\rangle$ and $\left|aa\right\rangle$ are oscillations at short timescale and finally reach to a steady state in the mixed phase, which are different from that in PI. The different populations evolution corresponds to the distinctive entanglement dynamics in the mixed phase and PI. Notably, the populations of basis state are equal at specific time ${{T}^{*}}=27.2$, i.e., $\left|\alpha\right|\sim\left|\beta\right|\sim\left|\varsigma\right|\sim\left|\delta\right|$.
Figure C3: The decay rate of the upper (${{\gamma}_{up}}$) and lower (${{\gamma}_{low}}$) envelopes in the weak (a-c) and strong (d-f) coupling regimes. The dashed lines and solid lines are the fitting envelopes for the type III entanglement dynamics and concurrence evolution.
Furthermore, we compare the decay rate of the fitting envelope in different coupling strength. We can find that both the decay rate of the upper (${{\gamma}_{up}}$) and lower envelopes (${{\gamma}_{low}}$) are decrease with the increase of the coupling strength. Additionally, in the weak coupling regime, the decay rates of the upper envelope are larger than that of the lower envelope, i.e., ${{\gamma}_{up}}>{{\gamma}_{low}}$ (see Fig. C3 (a-c)). On the contrary, the decay rates of the upper envelope are smaller than that of the lower envelope, i.e., ${{\gamma}_{up}}<{{\gamma}_{low}}$, in the strong coupling regime (see Fig. C3 (d-f)). Additionally, we can only get the lower envelope fitting in the region of PI, and the relationship between the decay rate of upper and lower envelope fittings are not obvious.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) (Grants No. 62075004, No. 11804018, No. U23A20481, No. 62275010, No. 12474353, and No. 12474354), Beijing Municipal Natural Science Foundation (Grants No. 4212051 and No. 1232027), and the Fundamental Research Funds for the Central Universities.